OpenStudy (toribyrd33):

Joselyn is a manager at a sign painting company. She has three painters, Allen, Brianne, and Charles. Allen can complete a large project in 16 hours. Brianne can complete the same sized project in 18 hours. Charles is new, so no one knows how long it will take him. Joselyn assigns them all a large project to complete together. Explain to Joselyn how this project can tell her how long it would take Charles if he worked by himself. Use complete sentences.

OpenStudy (dtan5457):

Would it not just be that if Charles works alone, we can get the amount of hours he can do a large project by himself? just like Allen=16 hours and Brianne=18 hours

OpenStudy (toribyrd33):

I guess but they don't give any indication as to how long it takes Charles to do anything

OpenStudy (mimi_x3):

I dont get this question tbh

OpenStudy (mimi_x3):

THIS IS NOT EVEN MATH!!!!!!!!!!

OpenStudy (dan815):

okay listen

OpenStudy (mimi_x3):

dann shhhh

OpenStudy (dtan5457):

With out a single piece of information about Charles, I don't know how anything can solved

OpenStudy (dan815):

A completes some work in 18 hrs B completes it in 16 hrs X completes it in N hrs

OpenStudy (dan815):

they are all working on the same work W

OpenStudy (mimi_x3):

If they all worked together how long wld it take???

OpenStudy (dan815):

rate of A =W/18 rate of B=W/16 rate of X= W/n therefore let t be the time they complete work (W/18+w/16+w/n)*t=W solve for n

OpenStudy (dan815):

u can notice that w is actually common and independant in this case

OpenStudy (dan815):

(W/18+w/16+w/n)*t=W = (1/18+1/16+1/n)*t=1

OpenStudy (dan815):

make sense?

OpenStudy (dan815):

u can think more intuively like um

OpenStudy (dan815):

say i walk 2 steps every hour

OpenStudy (dan815):

no lets change that

OpenStudy (dan815):

say i walk 2 steps every second, suppose rachel walks 3 steps every second

OpenStudy (dan815):

and there is a "work" which requires 50 steps to be walked

OpenStudy (dan815):

2*t+3*t=50 t=10, in 10 secs id have walked 20 and rachel walked 30 to give us 50 total steps

OpenStudy (mimi_x3):

not realistic whatsoever -.-

OpenStudy (dan815):

another way to think of it is

OpenStudy (toribyrd33):

wait going back to the equation would you just solve for x?

OpenStudy (dan815):

50/2=25 secs I alone complete it in 25 secs rachel alone completes it in 50/3 seconds

OpenStudy (dan815):

u cannot solve for it that is an expression there are 1 too many unknowns in this case

OpenStudy (dan815):

u are simply explaining how given you know the time it takes for all 3 of them to complete the work u can tell the rate at which x works

OpenStudy (toribyrd33):

okay so we're not actually finding out how long it takes Charles

OpenStudy (mimi_x3):

We are giving a method of how to solve if we were perchance given some extra info

OpenStudy (toribyrd33):

oh ok I think i know how to do that